### v2

@@ 4 Chronology violating pseudo-timelike loops
In Section 2, we introduced the notion of closed pseudo-timelike curves on a signature-type changing background and we demonstrated how they must be defined to ensure that the concept of causality remains meaningful. In this section, we will reveal the non-well-behaved nature of transverse, signature-type changing, nn–dimensional manifolds with a transverse radical.
@@ 4.1 Local pseudo-timelike loops
Clearly, $\tilde{g}(X,X)\geq 0$ because $G_{0}>0$ per definition and $\delta_{jk}X^{j}X^{k}=\frac{t(X^{0})^{2}}{G_{0}}\geq 0$. Therefore, the vector $X\in T_{p}M,\>p\in C_{0}$ is not timelike with respect to $\tilde{g}$. This means, within $C_{0}$ the $\tilde{g}$–light cones always reside inside of the $\tilde{g}_{0}$–light cones, i.e. $\tilde{g}\leq\tilde{g}_{0}$ in $C_{0}$. The cull cones of $\tilde{g}_{0}$ are more opened out than those of the metric $\tilde{g}$. Denote $p_{0}\in C_{0}$ by $(t(p_{0}),x^{1}(p_{0}),\ldots,x^{n-1}(p_{0}))=(t_{0},x_{0}^{1},\ldots,x_{0}^{n-1})$.
@@
Since $\sigma$ is given as a function of $t$, it represents the arc length from the starting point at $t(p_{0})=t_{0}$ to $t(0)=0$. Then past-directed $\tilde{g}_{0}$–lightlike curves emanating from $p_{0}$ reach the hypersurface at $t=0$ after passing through the arc length distance
@@
Provided this arc length distance satisfies $\triangle\sigma\leq\xi_{0}$, then the past-directed lightlike curves $\alpha(t)$ (emanating from $p_{0}$) reach the hypersurface at $t=0$ while remaining within $C_{0}$. Accordingly this is also the case for $\tilde{g}$–lightlike curves emanating from $p_{0}$. Conversely, if $\triangle\sigma>\xi_{0}$ then there exist past-directed $\tilde{g}_{0}$–lightlike curves emanating from $p_{0}$ that reach the hypersurface outside of $C_{0}$.
In this case, we have $\triangle\sigma=\frac{2}{3}\sqrt{\frac{t_{0}^{3}}{G_{0}}}>\xi_{0}\Longleftrightarrow t_{0}>\sqrt[3]{\frac{9}{4}\xi_{0}^{2}\cdot G_{0}}$ and we must adjust the new starting point $p_{1}=(t_{1},x_{0})$ accordingly by setting $t_{1}\leq\sqrt[3]{\frac{9}{4}\xi_{0}^{2}\cdot G_{0}}<t_{0}$. Thereby we make sure that all past-directed $\tilde{g}$–lightlike curves emanating from $p_{1}$ hit the hypersurface $\mathcal{H}$ without leaving $C_{0}$. That is $\overline{I_{\tilde{g}}^{-}(p_{1})}\subset C_{0}\subset U_{\varphi}\subset M$, where $\overline{I_{\tilde{g}}^{-}(p_{1})}$ is the $\tilde{g}$–chronological past of the event $p_{1}\in M_{L}$, restricted to $M_{L}\cup\mathcal{H}$, see Figure 6.
@@
The above corollary follows directly from Theorem 2 because we have proven that all past-directed $\tilde{g}$lightlike curves emanating from $p_{1}$ hit the hypersurface $\mathcal{H}$ without leaving $C_{0}$. That is $\overline{I_{\tilde{g}}^{-}(p_{1})}\subset C_{0}\subset U_{\varphi}\subset M$, where $\overline{I_{\tilde{g}}^{-}(p_{1})}$ is the closure of the $\tilde{g}$–chronological past of the event $p_{1}\in M_{L}$. Hence, we have also shown that the causal past is within $C_{0}$, and furthermore, $J_{\tilde{g}}^{-}(p_{0})\subset\overline{I_{\tilde{g}}^{-}(p_{0})}$.
@@
We note that the loops constructed in the proof of Theorem 2 are entirely contained within the small neighborhoods $U(q)$ and do not extend globally. More generally, pseudo-timelike loops may exist that traverse larger portions of the manifold, particularly if the hypersurface $\mathcal{H}$ has a toroidal or otherwise nontrivial global structure. Allowing self-intersections does not conflict with treating the objects as parameterized curves: by definition, loops are parameterized, unlike general closed curves, so the parametrization is fixed from the start.
@@ 4.2 Global pseudo-timelike loops
The existence of such pseudo-timelike curves locally near the hypersurface that loop back to themselves, gives rise to the question whether this type of curves also occur globally. We want to elucidate this question in the following. [^21] A spacetime is a Lorentzian manifold that models space and time in general relativity and physics. This is conventionally formalized by saying that a spacetime is a smooth connected time-orientable Lorentzian manifold $(M,g)$ with $\dim M=4$. But in what follows we want to study the nn–dimensional ($n\geq 2$) case.
@@
Due to [4] we know that $M_{L}$ admits a splitting $M_{L}=(\mathbb{R}_{>0})_{t}\times S_{t}=\bigcup_{t\in\mathcal{\mathbb{R}}_{>0}}S_{t}$, such that the Lorentzian sector $M_{L}$ is decomposed into hypersurfaces (of dimension n1n-1), specified as the level surfaces $S_{t}=\mathcal{T}^{-1}(t)=\{p\in M_{L}\colon\mathcal{T}(p)=t\},t\in\mathbb{R}_{>0}$, of the real-valued smooth temporal function $\mathcal{T}\colon M_{L}\longrightarrow\mathbb{R}_{>0}$ whose gradient $\textrm{grad}\mathcal{T}$ is everywhere non-zero and, clearly, $d\mathcal{T}$ is an exact 11–form. Within the neighborhood $U=\bigcup_{q\in\mathcal{H}}U(q)$ this foliation $\bigcup_{t\in\mathcal{\mathbb{R}}_{>0}}S_{t}$ can be chosen in such a way that it agrees with the natural foliation given by the absolute time function $\mathfrak{h}(t,\mathbf{\hat{x}}):=t$, see Remark 6 and Definition 16.242424Recall that a smooth function $T\colon M\longrightarrow\mathbb{R}$ on a connected time-orientable Lorentzian manifold $(M,g)$ is a global time function if $T$ is strictly increasing along each future-pointing non-spacelike curve. Moreover, a temporal function is a time function $T$ with a timelike gradient $\textrm{grad}T$ everywhere. Since $M_{L}$ is globally hyperbolic it admits a smooth global time function $T$ and consequently it admits [31] a temporal function $\mathcal{T}$. Hence, in the Lorentzian sector $M_{L}$ there exists a global temporal function $\mathcal{T}\colon M_{L}\longrightarrow\mathbb{R}_{>0}$, and $\textrm{grad}\mathcal{T}$ is orthogonal to each of the level surfaces $S_{t}=\mathcal{T}^{-1}(t)=\{p\in M_{L}\colon\mathcal{T}(p)=t\},t\in\mathbb{R}_{>0}$, of $\mathcal{T}$. Note that $\mathcal{T}=t$ is a scalar field on $M_{L}$, hence $\textrm{grad}\mathcal{T}=\textrm{grad}t=(dt)^{\#}$.
@@
Theorem 3 explicitly states that through every point in $M$, there always exists a pseudo-timelike loop. Consequently, this applies equally to points on the hypersurface as well as to those in the Riemannian region. In such cases, the construction necessarily produces a pseudo-timelike loop with two intersection points. Indeed, the point $p$ is required to serve as a point of self-intersection. Consequently, the pseudo-timelike loop must also admit an intersection in $M_{L}$, and at least one of the intersection points must lie in $M_{L}$ in order for the curve to qualify as a pseudo-timelike loop. The second intersection point is then determined, either as $p$ on $\mathcal{H}$ or as $p$ in $M_{R}$. Topologically, a loop with two intersection points falls within the class of curves described in Definition 9.
@@ Acknowledgments
NER is greatly indebted to Richard Schoen for generously welcoming her into his research group and to Alberto Cattaneo for affording her with creative independence throughout the duration of this research endeavor. Moreover, NER acknowledges partial support of the SNF Grant No. 200021-227719. This research was (partly) supported by the NCCR SwissMAP, funded by the Swiss National Science Foundation.

### v3

@@ 4 Chronology violating pseudo-timelike loops
In Section 2, we introduced the notion of closed pseudo-timelike curves on a signature-type changing background and we demonstrated how they must be defined to ensure that the concept of causality remains meaningful. In this section, we will reveal the non-well-behaved nature of transverse, signature-type changing, $n$-dimensional manifolds with a transverse radical.
@@ 4.1 Local pseudo-timelike loops
Clearly, $\tilde{g}(X,X)\geq 0$ because $G_{0}>0$ per definition and $\delta_{jk}X^{j}X^{k}=\frac{t(X^{0})^{2}}{G_{0}}\geq 0$. Therefore, the vector $X\in T_{p}M,\>p\in C_{0}$ is not timelike with respect to $\tilde{g}$. This means, within $C_{0}$ the $\tilde{g}-$light cones always reside inside of the $\tilde{g}_{0}$-light cones, i.e. $\tilde{g}\leq\tilde{g}_{0}$ in $C_{0}$. The cull cones of $\tilde{g}_{0}$ are more opened out than those of the metric $\tilde{g}$. Denote $p_{0}\in C_{0}$ by $(t(p_{0}),x^{1}(p_{0}),\ldots,x^{n-1}(p_{0}))=(t_{0},x_{0}^{1},\ldots,x_{0}^{n-1})$.
@@
Since $\sigma$ is given as a function of $t$, it represents the arc length from the starting point at $t(p_{0})=t_{0}$ to $t(0)=0$. Then past-directed $\tilde{g}_{0}$-lightlike curves emanating from $p_{0}$ reach the hypersurface at $t=0$ after passing through the arc length distance
@@
Provided this arc length distance satisfies $\triangle\sigma\leq\xi_{0}$, then the past-directed lightlike curves $\alpha(t)$ (emanating from $p_{0}$) reach the hypersurface at $t=0$ while remaining within $C_{0}$. Accordingly this is also the case for $\tilde{g}-$lightlike curves emanating from $p_{0}$. Conversely, if $\triangle\sigma>\xi_{0}$ then there exist past-directed $\tilde{g}_{0}$-lightlike curves emanating from $p_{0}$ that reach the hypersurface outside of $C_{0}$.
In this case, we have $\triangle\sigma=\frac{2}{3}\sqrt{\frac{t_{0}^{3}}{G_{0}}}>\xi_{0}\Longleftrightarrow t_{0}>\sqrt[3]{\frac{9}{4}\xi_{0}^{2}\cdot G_{0}}$ and we must adjust the new starting point $p_{1}=(t_{1},x_{0})$ accordingly by setting $t_{1}\leq\sqrt[3]{\frac{9}{4}\xi_{0}^{2}\cdot G_{0}}<t_{0}$. Thereby we make sure that all past-directed $\tilde{g}-$lightlike curves emanating from $p_{1}$ hit the hypersurface $\mathcal{H}$ without leaving $C_{0}$. That is $\overline{I_{\tilde{g}}^{-}(p_{1})}\subset C_{0}\subset U_{\varphi}\subset M$, where $\overline{I_{\tilde{g}}^{-}(p_{1})}$ is the $\tilde{g}-$chronological past of the event $p_{1}\in M_{L}$, restricted to $M_{L}\cup\mathcal{H}$, see Figure 6.
@@
The above corollary follows directly from Theorem 2 because we have proven that all past-directed $\tilde{g}-$lightlike curves emanating from $p_{1}$ hit the hypersurface $\mathcal{H}$ without leaving $C_{0}$. That is $\overline{I_{\tilde{g}}^{-}(p_{1})}\subset C_{0}\subset U_{\varphi}\subset M$, where $\overline{I_{\tilde{g}}^{-}(p_{1})}$ is the closure of the $\tilde{g}-$chronological past of the event $p_{1}\in M_{L}$. Hence, we have also shown that the causal past is within $C_{0}$, and furthermore, $J_{\tilde{g}}^{-}(p_{0})\subset\overline{I_{\tilde{g}}^{-}(p_{0})}$.
@@
We note that the loops constructed in the proof of Theorem 2 are entirely contained within the small neighborhoods $U(q)$ and do not extend globally. More generally, pseudo-timelike loops may exist that traverse larger portions of the manifold, particularly if the hypersurface $\mathcal{H}$ has a toroidal or otherwise nontrivial global structure. Allowing self-intersections does not conflict with treating the objects as parameterized curves.
@@ 4.2 Global pseudo-timelike loops
The existence of such pseudo-timelike curves locally near the hypersurface that loop back to themselves, gives rise to the question whether this type of curves also occur globally. We want to elucidate this question in the following. [^21] A spacetime is a Lorentzian manifold that models space and time in general relativity and physics. This is conventionally formalized by saying that a spacetime is a smooth connected time-orientable Lorentzian manifold $(M,g)$ with $\dim M=4$. But in what follows we want to study the $n$-dimensional ($n\geq 2$) case.
@@
Due to [4] we know that $M_{L}$ admits a splitting $M_{L}=(\mathbb{R}_{>0})_{t}\times S_{t}=\bigcup_{t\in\mathcal{\mathbb{R}}_{>0}}S_{t}$, such that the Lorentzian sector $M_{L}$ is decomposed into hypersurfaces (of dimension n1n-1), specified as the level surfaces $S_{t}=\mathcal{T}^{-1}(t)=\{p\in M_{L}\colon\mathcal{T}(p)=t\},t\in\mathbb{R}_{>0}$, of the real-valued smooth temporal function $\mathcal{T}\colon M_{L}\longrightarrow\mathbb{R}_{>0}$ whose gradient $\textrm{grad}\mathcal{T}$ is everywhere non-zero and, clearly, $d\mathcal{T}$ is an exact $1$-form. Within the neighborhood $U=\bigcup_{q\in\mathcal{H}}U(q)$ this foliation $\bigcup_{t\in\mathcal{\mathbb{R}}_{>0}}S_{t}$ can be chosen in such a way that it agrees with the natural foliation given by the absolute time function $\mathfrak{h}(t,\mathbf{\hat{x}}):=t$, see Remark 6 and Definition 16.242424Recall that a smooth function $T\colon M\longrightarrow\mathbb{R}$ on a connected time-orientable Lorentzian manifold $(M,g)$ is a global time function if $T$ is strictly increasing along each future-pointing non-spacelike curve. Moreover, a temporal function is a time function $T$ with a timelike gradient $\textrm{grad}T$ everywhere. Since $M_{L}$ is globally hyperbolic it admits a smooth global time function $T$ and consequently it admits [31] a temporal function $\mathcal{T}$. Hence, in the Lorentzian sector $M_{L}$ there exists a global temporal function $\mathcal{T}\colon M_{L}\longrightarrow\mathbb{R}_{>0}$, and $\textrm{grad}\mathcal{T}$ is orthogonal to each of the level surfaces $S_{t}=\mathcal{T}^{-1}(t)=\{p\in M_{L}\colon\mathcal{T}(p)=t\},t\in\mathbb{R}_{>0}$, of $\mathcal{T}$. Note that $\mathcal{T}=t$ is a scalar field on $M_{L}$, hence $\textrm{grad}\mathcal{T}=\textrm{grad}t=(dt)^{\#}$.
@@
Theorem 3 explicitly states that through every point in $M$, there always exists a pseudo-timelike loop. Consequently, this applies equally to points on the hypersurface as well as to those in the Riemannian region. In such cases, the construction necessarily produces a pseudo-timelike loop with two intersection points. Indeed, the point $p$ is required to serve as a point of self-intersection. Consequently, the pseudo-timelike loop must also admit an intersection in $M_{L}$, and at least one of the intersection points must lie in $M_{L}$ in order for the curve to qualify as a pseudo-timelike loop. The second intersection point is then determined, either as $p$ on $\mathcal{H}$ or as $p$ in $M_{R}$. [^28] The construction for points $p\in\mathcal{H}$ necessitates that the timelike curve segments from $M_{L}$ approach the “pinched” light cones near the hypersurface. Consequently, the curve segments that pass through $\mathcal{H}$ into the Riemannian sector $M_{R}$ and self-intersect at $p\in\mathcal{H}$ must meet with first-order contact. Topologically, a loop with two intersection points falls within the class of curves described in Definition 9.
@@ Acknowledgments
The authors thank the reviewers for their insightful comments and critique, which have helped to refine and improve this article. NER is greatly indebted to Richard Schoen for generously welcoming her into his research group and to Alberto Cattaneo for affording her with creative independence throughout the duration of this research endeavor. Moreover, NER acknowledges partial support of the SNF Grant No. 200021-227719. This research was (partly) supported by the NCCR SwissMAP, funded by the Swiss National Science Foundation.